OpenStudy (anonymous):

Izzy's heartbeat rate was measured on an EKG machine. After conversion, the function produced could be modeled by a sine function, and the wave produced a maximum of 8, minimum of −2, and period of π. Which of the following functions could represent Izzy's EKG read-out?

OpenStudy (deadly_roses):

f(x) = 5 sin 2x + 3 should be the answer when done...since all other ways cant match the read-out properly....

OpenStudy (ybarrap):

 We need amplitude, A and period, T: $$ A\sin\frac{2\pi t}{T} $$ The amplitude goes from 8 to -2 so the max "swing" is 8 - (-2) = 10, the amplitude is 10/2 = 5 = A, half the "swing", and the period, T is \(\pi\). Does this make sense?